OpenStudy (anonymous):

a rectangular prism has a length x+2 width x+1 heith 4 and volume 24 find length and width WILL MEDAL PLEASE HELP

OpenStudy (lieutenantgeneral):

Volume = height x width x length V = 24 = 4 (x+2)(x+1)....................so 6 = (x+2)(x+1) x^2 + 3x + 2 - 6 = 0 x^2 + 3x - 4 = 0 ( x + 4)(x - 1) = 0 x = -4 or x =1........disregard negative number x = 1 So length = 1+2 = 3.............. and width = 2 Check volume = 3*2*4 = 24.................correct